OpenStudy (anonymous):

how do you do this problem: (a+b)x^2+ (a+b)x-20(a+B)

OpenStudy (anonymous):

assuming you want to factor this: = (a+b)(x^2 + x - 20) = (a+b)(x+5)(x-4)

OpenStudy (anonymous):

Thanks Jimmy rep, I need that to figure out how to reverse. :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

on this assignment I need to reverse on this problem, so I kind of got lost what formula.